# VascX Models: Model Ensembles for Retinal Vascular Analysis from Color Fundus Images

### v1

@@ Abstract
We introduce VascX models, a comprehensive set of model ensembles for analyzing retinal vasculature from color fundus images (CFIs). Annotated CFIs were aggregated from public datasets for vessel, artery-vein, and disc segmentation; and fovea localization. Additional CFIs from the population-based Rotterdam Study were, with arteries and veins annotated by graders at pixel level. Our models achieved robust performance across devices from different vendors, varying levels of image quality levels, and diverse pathologies. Our models demonstrated superior segmentation performance compared to existing systems under a variety of conditions. Significant enhancements were observed in artery-vein and disc segmentation performance, particularly in segmentations of these structures on CFIs of intermediate quality, a common characteristic of large cohorts and clinical datasets. Our model outperformed human graders in segmenting vessels with greater precision. With VascX models we provide a robust, ready-to-use set of model ensembles and inference code aimed at simplifying the implementation and enhancing the quality of automated retinal vasculature analyses. The precise vessel parameters generated by the model can serve as starting points for the identification of disease patterns in and outside of the eye.
@@ Introduction
Today the understanding of retinal vasculature and its association with diseases has expanded substantially. A significant focus has been placed on biomarkers such as retinal calibers, tortuosity, and branching characteristics. These biomarkers are now studied not only in relation to ocular diseases but also to systemic conditions, particularly cardiovascular diseases [37, 24]. More recently, attention has turned to the retina as a window into brain pathology, with studies linking retinal vascular changes to neurodegenerative diseases such as Alzheimer’s disease [12].
Parallel to the scientific advances, deep learning brought about ground-breaking improvements in segmentation quality across a variety of tasks [19]. This triggered the development of multiple datasets and deep learning architectures for the segmentation of the main anatomical structures visible in CFIs: vessel segmentation, artery-vein segmentation, and disc segmentation [TODO]. More recent are attempts to integrate these models into pipelines for fully automated vascular feature extraction. Automorph is the most notable example of a deep-learning-based pipeline integrating vessel segmentation, artery-vein segmentation, disc segmentation, and vascular feature extraction [40]. Additionally, region-based analysis have been implemented to isolate features from specific areas of the retina, utilizing optic disc and fovea detection models. These approaches allow for a more granular analysis, offering a deeper understanding of the temporal development of retinal changes.
Despite these developments, automated segmentation of features on CFIs remains far from solved. Publicly available training datasets are small and homogeneous, often limited to images from a single device, centered on small anatomic regions like the fovea or optic disc, or with a narrow field of view. Until recently [35], the largest publicly available dataset [2] for artery-vein segmentation contained only 45 images. Additionally, the pathology and demographics of the study population are often restricted, which significantly impacts the generalizability of the models trained on these datasets. These limitations are particularly problematic in analyses of large cohort studies, which typically encompass a diverse range of devices and capture conditions. To overcome these drawbacks, large and diverse datasets are needed that allow the development of robust models that are effective and reliable across a broad spectrum of conditions Annotations of vessel crossings should be optimized to ensure modeling of intact vessel tree connectivity.
@@
  • We present models for optic disc, vessel, and artery-vein segmentation and fovea localization, fundamental models in the development of CFI analysis systems, especially of the retinal vasculature. We built our models using public datasets and new annotations by professional graders on diverse sets of CFIs from the Rotterdam Study.
@@ 0.1 Datasets
We trained and evaluated our models on a combination of public datasets and images from the Rotterdam Study (RS).
0.1.1 Public Datasets
@@
We collected publicly available datasets from previous work. These datasets have different patient demographics, pathologies, retinal regions (fovea and disc centered), capture devices, field of view and countries of origin. Table 1 shows an overview of these datasets.
@@
Table 1: Overview of datasets for various segmentation and localization tasks. Pathologies: H: healthy, GC - glaucoma (all types), NTG: normal tension glaucoma, HTG: high tension glaucoma, HT: hypertension, DR: diabetic retinopathy, AMD: age-related macular degeneration, PM: pathological myopia, O: other eye diseases. * 379 of 1200 samples were discarded due to containing no disc segmentation for partially visible discs. ** 21 of 1200 CFIs were discarded due to having no associated annotation.
@@
Dataset N Origin Ages Pathology Region FOV Vessel Segmentation Chase DB [11] 28 UK Children - OD 35° DRHAGIS [16] 40 UK NR 10 GC, 10 HT, 10 DR, 10 AMD M 45° HRF [2] 45 GER NR 15 H, 15 DR, 15 GC M 45° RETA [25] 54 IND NR Signs of DR M 50° FIVES [21] 800 CN 4-83 H, GC, DR, AMD M 50° Leuven-Haifa [35] 240 BE 18-90 75 NTG, 63 HTG, 56H, 30 O OD 30° RS (ours) 352 NL 40+ See Figure 1. Total 1559 Artery Vein Segmentation RITE [31] 40 NL 25-90 7 DR, 33 no DR M 45° HRF-AV [3] 45 GER NR 15 H, 15 DR, 15 GC M 45° Les-AV [26] 22 NR NR NR OD 30° Leuven-Haifa [35] 240 BE 18-90 75 NTG, 63 HTG, 56H, 30 O OD 30° RS (ours) 215 NL 40+ See Figure 1. Total 562 Disc Segmentation ORIGA [39] 650 SGP 40-80 Multiple OD NR PAPILA [23] 488 SPN 15-90 87 GC, 333 H, 68 O OD 30 ° IDRiD [30] 81 IND NR Signs of DR M 50° ADAM [8] 821* CN $53.2\pm 15.6$ 267 AMD, 933 O M, OD, MP 45 ° PALM [7] 1179** CN $37.5\pm 15.91$ 637 PM, 563 O M, OD, MP 45 ° REFUGE2 [9] 2000 CN NR 280 GC, 1720 H/O M, OD, MP 45 ° RS (ours) 1225 NL 40+ See Figure 1. Total 7464 Fovea Localization IDRiD [30] 516 IND NR Signs of DR M 50° ADAM [8] 1200 CN $53.2\pm 15.6$ 267 AMD, 933 O M, OD, MP 45 ° PALM [7] 1200 CN $37.5\pm 15.91$ 637 PM, 563 O M, OD, MP 45 ° REFUGE2 [9] 2000 CN NR 280 GC, 1720 H/O M, OD, MP 45 ° RS (ours) 10908 NL 40+ See Figure 1. Total 15824 CFI Quality Estimation EyeQ [14] 28,792
@@
We included several well-known vessel and artery-vein segmentation datasets with diabetic retinopathy (DR) patients [16, 2, 25, 3]. We excluded the DRIVE dataset [32, 31] due to its low resolution ($584\times 565$ pixels). We included the larger and more recent FIVES [21] and Leuven-Haifa [35] datasets, which contain a mixture of diseased and healthy patients. On the other hand, most disc segmentation datasets were originally collected for the assessment of glaucoma [39, 23, 9]. Exceptions are ADAM (age-related macular degeneration), PALM (pathological myopia) and IDRiD (diabetic retinopathy); which also contain fovea positions. Finally, we used the Eye-Quality (EyeQ) dataset for the quality estimation algorithm. This dataset is unique for its size and for containing a variety of capture devices and image conditions. The field-of-view of the images in our training set is between 30°and 50°, which includes the vast majority of standard fundus imaging devices.
0.1.2 Rotterdam Study Datasets
@@
We complimented the CFI’s from the public data sets with randomly sampled CFIs from the Rotterdam Study (RS). This study is a prospective population-based cohort study of people living in Ommoord, a district of the city of Rotterdam ( [18]). The RS consists of four cohorts, all of which were used in this work. The minimum age of study participants varies between $>55$ in the first cohort and $>40$ in the fourth. Each cohort was followed for multiple rounds of follow-up examinations every 4 to 5 years. Most of the patient visits in the RS involved the capture of CFIs on both eyes. Due to the multi-decade span of the RS, multiple devices, capture conditions and fields (macula and disc centered) are present in the dataset. The RS is entered into the Netherlands National Trial Register (NTR) and the WHO International Clinical Trials Registry Platform (ICTRP) under shared catalog number NTR6831. All participants provided written informed consent following the Declaration of Helsinki to participate in the study and to have their information obtained from their treating physicians.
0.1.3 CFI Selection and Weighing
@@
We used most public datasets in their entirety without filtering, after harmonizing annotations where necessary. The main exception were disc segmentations in the ADAM dataset, where 379 of 1200 samples were discarded due to containing no disc segmentation for partially visible discs.
@@
Regarding the Rotterdam Study, we selected CFIs at random for the vessel and A/V segmentation datasets (Table 1), only excluding images that were deemed unusable due to low image quality. We used the quality estimation model in Automorph [40] to assess image quality. Additionally, we asked graders to exclude images for which they found the quality to be unusable for the task at hand.
@@
For disc segmentation, due to the large size of publicly available datasets (Table 1), we used a model trained on publicly available datasets to automatically identify failure or challenging CFIs. We selected CFIs for which no disc was detected or a disconnected disc mask was produced by the model. A trained grader reviewed these images and identified the ones where the disc was present but incorrectly segmented, and these were included in the training set.
@@
For fovea detection, we used CFIs from a RS sub-cohort graded for Age-related Macular Degeneration (AMD).
@@
Figure 1 shows the statistics of the final Rotterdam Study CFI development sets. Vessel and artery-vein segmentation sets are similar (partly due to overlap) and are closer to the overall distribution of the RS. The disc segmentation dataset, on the other hand, contains more images from the analog devices, which were often the hardest to segment partly due to unconventional disc placement.
@@
Figure 1: Statistics of the Rotterdam Study sets used to train VascX models.
0.2 CFI annotation
@@
Images were annotated using custom software for ophthalmological image annotation. Four professional graders participated in the process. The graders worked on desktop computers with a drawing tablet for segmentation. The annotation process was different per annotation type:
@@
For the first 100 images the interface was initialized with outputs from the publicly available model Automorph [40]. The graders were instructed to independently fix the masks produced by the model, including missing vessels and over-segmentation. A digitally enhanced version of the image with increased contrast of vessel edges was used to aid the graders (Figure 2(a)). After this, a model trained on these annotations and the publicly available data was used in place of Automorph as initialization for the remaining images.
@@
We resolved connectivity issues at artery-vein crossings by annotating arteries and veins on separate layers. An interface was developed specifically for artery-vein segmentation (Figure 2(b)) to allow the graders to start from a vessel segmentation and color arteries and veins into two independent masks or layers using drawing tools. A third layer - Unknown was added for vessels that could not be recognized as either. They were asked to color A/V crossings on both layers (overlapping). They were able to visualize and correct each layer independently, or all the same time. Another tool allowed them to visualize the connected components of each mask in different colors to easily find mistakes in connectivity. The entire process consisted in 1) correcting mistakes in the provided mask (vessel segmentation); 2) coloring the mask into A/V and unknown colors/layers; 3) Verifying connected components on the A/V masks and filling any gaps.
@@
Four graders placed an ETDRS grid on the fundus image. The process consisted in clicking on the fovea center and on the optic disc border point closest to the fovea. These points were used to display a grid on top of the image, which could be corrected in case of mistakes.
@@
Figure 2: Screen capture of the software used for all the annotations on Rotterdam Study CFIs.
0.3 Model Training
@@ 0.3.1 Data pre-processing
    Detection of the CFI boundaries defined by a circle and top and bottom vertical padding.
@@
The geometric steps above (cropping and resizing) were replicated on the labels for training.
0.3.2 Vessel and A/V segmentation models
@@
Vessel annotations were encoded as 0) background and 1) vessel. Artery-vein annotations were encoded as 0) background, 1) arteries, 2) veins, and 3) crossings (both artery and vein). We used this approach with the goal of boosting the relative importance of the crossings in the loss function. The loss function was the weighted sum of Dice and Cross-entropy with equal weights.
@@
0.3.3 Fovea and disc localization models
The fovea and disc localization models were trained separately to localize keypoints on the CFIs. A heatmap regression approach was used using the same U-Net architecture used for artery-vein segmentation. Annotated keypoints were used to generate target heatmaps ($512x512px$) using a Gaussian centered at the keypoint location. The $\sigma$ for the Gaussian was set via experiments on a different CFI localization task. Mean squared error (MSE) was used as loss function to train the model to approximate the continuous target heatmaps. The location (2D-index) of the maximum over the heatmaps output by the model was taken as the keypoint location.
Input CFIs were augmented via random appearance transforms: de-focus, hue-saturation-value, Gamma and Gaussian noise; followed by geometric transforms: horizontal flipping, random scale (1.0-1.05 times original scale) random rotation (up to 5 degrees) and finally resizing down to model input size. Test-time augmentation was correspondingly applied using the vertical flip of the input.
0.3.4 Quality classification model
@@ 0.4 Evaluation
0.4.1 Quality evaluation
We made use of the Eye-Quality (EyeQ) Assessment Dataset [14] to train our quality evaluation model, which we used to assess the quality of the CFIs (Table 1). This dataset contains manual gradings of fundus quality using a 3-class system. CFIs were graded as one of: unusable/bad, usable and good quality. We took advantage of the ordinal nature of these classes to produce continuous quality scores from the outputs of our quality classification model:
  1. 1.
    Applied the Softmax function with a temperature parameter $T$ to obtain probabilities:
    $p_{\_}i=\frac{e^{z_{\_}i/T}}{\sum_{\_}{j=1}^{n}e^{z_{\_}j/T}}\quad\text{for}\quad i=1,2,\ldots,n$ (1)
    where $z_{\_}i$ is the logit corresponding to class $i$ and $n=3$; the number of classes.
  2. 2.
    Merged probabilities into a score by linearly combining the class indices and the probabilities:
    $S=\sum_{\_}{i=1}^{n}i\cdot p_{\_}i$ (2)
This approach takes into account all logits to produce the final score. The value of T influences the sharpness of the probabilities and therefore how much the quality score is dominated by the largest logit. We tuned $T$ manually to balance the sharpness and smoothness of the probability distribution of the resulting scores.

### v2

@@ Abstract
We introduce VascX models, a comprehensive set of model ensembles for analyzing retinal vasculature from color fundus images (CFIs). Annotated CFIs were aggregated from public datasets . Additional CFIs, mainly from the population-based Rotterdam Study were annotated by graders for arteries and veins at pixel level, resulting in a dataset diverse in patient demographics and imaging conditions. VascX models demonstrated superior segmentation performance across datasets, image quality levels, and anatomic regions when compared to existing, publicly available models, likely due to the increased size and variety of our training set. Important improvements were observed in artery-vein and disc segmentation performance, particularly in segmentations of these structures on CFIs of intermediate quality, common in large cohorts and clinical datasets. Importantly, these improvements translated into significantly more accurate vascular features when we compared features extracted from VascX segmentation masks with features extracted from segmentation masks generated by previous models. With VascX models we provide a robust, ready-to-use set of model ensembles and inference code aimed at simplifying the implementation and enhancing the quality of automated retinal vasculature analyses. The precise vessel parameters generated by the model can serve as starting points for the identification of disease patterns in and outside of the eye.
@@ Introduction
Today the understanding of retinal vasculature and its association with diseases has expanded substantially. A significant focus has been placed on biomarkers such as retinal calibers, tortuosity, and branching characteristics. These biomarkers are now studied not only in relation to ocular diseases but also to systemic conditions, particularly cardiovascular diseases [53, 34, 24]. More recently, attention has turned to the retina as a window into brain pathology, with studies linking retinal vascular changes to neurodegenerative diseases such as Alzheimer’s disease [17]. Previous work has shown that automatically extracted vascular features improve the performance of diabetic retinopathy (DR) detection and classification [2].
Parallel to the scientific advances, deep learning brought about ground-breaking improvements in segmentation quality across a variety of tasks [27]. This triggered the development of multiple datasets, deep learning architectures and models for the segmentation of the main anatomical structures visible in CFIs: vessel segmentation [32, 35], artery-vein segmentation [19, 15], and disc segmentation. The task of assessing fundus image quality has also received attention [18, 1, 21]. More recent are efforts to integrate these models into pipelines for fully automated, explainable, vascular feature extraction. Automorph is the most notable example of an open source deep-learning-based pipeline integrating vessel segmentation, artery-vein segmentation, disc segmentation, and vascular feature extraction [57]. Automorph has been applied in various studies including on the reproducibility of its features [50, 51, 20, 11]. The availability of disc segmentation models in Automorph allows for region-based analysis which isolate features from specific areas of the retina.
Despite these developments, the robust segmentation of retinal morphology (arteries, veins, optic disc, fovea) from CFIs across a variety of patient and imaging conditions remains unsolved. Most publicly available training datasets for artery-vein segmentation are small and homogeneous, often limited to images from a single device, centered on a single anatomic region (fovea or optic disc), and with a fixed field of view. Until recently [49], the largest publicly available dataset for artery-vein segmentation contained only 45 images [4]. Additionally, the pathology and demographics of the study population are often restricted, which significantly impacts the generalizability of the models trained on these datasets. This affects the downstream extraction of explainable vascular features, which depend on the quality of such segmentations [45, 11]. This is particularly problematic in analyses of large cohort studies and clinical data, which may encompass a diverse range of devices and capture conditions. To overcome these drawbacks, larger and more diverse training datasets are needed to develop models that work reliably across a broader spectrum of conditions.
@@
  • We present models for optic disc, vessel, and artery-vein segmentation and fovea localization, fundamental models in the development of CFI analysis systems, especially of the retinal vasculature. We built our models using public datasets and new annotations by professional graders on diverse sets of CFIs from Dutch studies, mainly the Rotterdam Study.
@@ Methods
We trained and evaluated our models on a combination of public datasets and images from the Rotterdam Study (RS) and other Dutch studies.
Public Datasets
@@
We collected publicly available datasets from previous work. These datasets have different patient demographics, pathologies, retinal regions (fovea and disc centered), field of view and countries of origin. Table 1 shows an overview of these datasets. For complimentary details including capture devices and resolution see Supplementary Table 4.
@@
Table 1: Overview of datasets for various segmentation and localization tasks. Pathologies: H: healthy, GC - glaucoma (all types), NTG: normal tension glaucoma, HTG: high tension glaucoma, HT: hypertension, DR: diabetic retinopathy, AMD: age-related macular degeneration, PM: pathological myopia, O: other eye diseases. *: 379 of 1200 samples were discarded due to containing no disc segmentation for partially visible discs. $\dagger$: 21 of 1200 CFIs were discarded due to having no associated annotation.
@@
Dataset N Origin Ages Pathology Region FOV Vessel Segmentation Chase DB [16] 28 UK Children - OD 35° DRHAGIS [23] 40 UK NR 10 GC, 10 HT, 10 DR, 10 AMD M 45° HRF [4] 45 GER NR 15 H, 15 DR, 15 GC M 45° RETA [36] 54 IND NR Signs of DR M 50° FIVES [29] 800 CN 4-83 H, GC, DR, AMD M 50° Leuven-Haifa [49] 240 BE 18-90 75 NTG, 63 HTG, 56H, 30 O OD 30° Rotterdam (ours) 352 NL 40+ See Figure 1. Total 1559 Artery Vein Segmentation RITE [43] 40 NL 25-90 7 DR, 33 no DR M 45° HRF-AV [6] 45 GER NR 15 H, 15 DR, 15 GC M 45° Les-AV [37] 22 NR NR NR OD 30° Leuven-Haifa [49] 240 BE 18-90 75 NTG, 63 HTG, 56H, 30 O OD 30° Rotterdam (ours) 215 NL 40+ See Figure 1. Total 562 Disc Segmentation ORIGA [56] 650 SGP 40-80 Multiple OD NR PAPILA [31] 488 SPN 15-90 87 GC, 333 H, 68 O OD 30 ° IDRiD [42] 81 IND NR Signs of DR M 50° ADAM [13] 821* CN $53.2\pm 15.6$ 267 AMD, 933 O M, OD, MP 45 ° PALM [12] 1179$\dagger$ CN $37.5\pm 15.91$ 637 PM, 563 O M, OD, MP 45 ° REFUGE2 [14] 2000 CN NR 280 GC, 1720 H/O M, OD, MP 45 ° Rotterdam (ours) 1225 NL 40+ See Figure 1. Total 7464 Fovea Localization IDRiD [42] 516 IND NR Signs of DR M 50° ADAM [13] 1200 CN $53.2\pm 15.6$ 267 AMD, 933 O M, OD, MP 45 ° PALM [12] 1200 CN $37.5\pm 15.91$ 637 PM, 563 O M, OD, MP 45 ° REFUGE2 [14] 2000 CN NR 280 GC, 1720 H/O M, OD, MP 45 ° Rotterdam (ours) 10908 NL 40+ See Figure 1. Total 15824 CFI Quality Estimation EyeQ [18] 28,792
@@
We included several well-known vessel and artery-vein segmentation datasets with diabetic retinopathy (DR) patients [23, 4, 36, 6]. We excluded the DRIVE dataset [46, 43] due to its low resolution ($584\times 565$ pixels). We included the larger and more recent FIVES [29] and Leuven-Haifa [49] datasets, which contain a mixture of diseased and healthy patients. On the other hand, most datasets with disc segmentations were originally collected for the assessment of glaucoma [56, 31, 14]. The REFUGE2 dataset, an extension of the original REFUGE dataset [38] is notable for its size (2000 CFIs), various anatomic regions captured, and for containing fovea location annotations. We also included ADAM (age-related macular degeneration) [13], PALM (pathological myopia) [12] and IDRiD (diabetic retinopathy) [42]; which also contain fovea positions.
Finally, we used the Eye-Quality (EyeQ) dataset (Table 1) for the quality estimation algorithm. This dataset contains manual gradings of fundus quality from 28,792 images using a 3-class system. CFIs were graded as one of: unusable/bad, usable and good quality. It is unique for its size and for containing a variety of capture devices and image conditions. The field-of-view of the images in our training set is between 30°and 50°, which includes the vast majority of standard fundus imaging devices.
Rotterdam Datasets
@@
We complimented the CFIs from the public data sets with images available to the Department of Ophthalmology at Erasmus Medical Center in Rotterdam; mainly the Rotterdam Study (RS). The RS is a prospective population-based cohort study of people living in Ommoord, a district of the city of Rotterdam ( [26]). The RS consists of four cohorts, all of which were used in this work. The minimum age of study participants varies between $>55$ in the first cohort and $>40$ in the fourth. Each cohort was followed for multiple rounds of follow-up examinations every 4 to 5 years. Most of the visits in the RS involved the capture of CFIs on both eyes. Due to the multi-decade span of the RS, multiple devices, capture conditions and fields (macula and disc centered) are present in the dataset.
In addition to the RS, we included images captured during the Randomized Controlled Trial for Age-Related Macular Degeneration (AMD-Life) [9]; which consisted of 150 patients aged 55–85 years imaged with a Topcon 3D OCT 2000-plus and Topcon TRC-50EX cameras. Images from the Dutch Myopia Study (MYST) [22] which included younger subjects (mean (SD) age of 47.3 (12.9) years) were also used. These images were captured with a Topcon TRC-50EX.
CFI Selection and Weighing
@@
We used most public datasets in their entirety without filtering, after harmonizing annotations when necessary. The main exception were disc segmentations in the ADAM dataset, where 379 of 1200 samples were discarded due to containing no disc segmentation for partially visible discs.
@@
Regarding the Rotterdam sets, we sampled CFIs at random from all available images. We included fundus images captured by OCT machines to improve diversity of devices in the dataset. For the vessel and A/V segmentation datasets (Table 1), due to the much larger size of the Rotterdam Study compared to AMD-Life and MYST, we over-sampled CFIs from these sets to make up 15% of the samples. We used the quality estimation model in Automorph [57] to assess image quality and excluded images classified as unusable. Additionally, we asked graders to exclude images for which they found the quality to be unusable for the task at hand.
@@
For disc segmentation, due to the large size of publicly available datasets (Table 1), we first used a model trained on publicly available datasets to automatically identify failure or challenging CFIs. We selected CFIs for which no disc was detected or a disconnected disc mask was produced by the model. A trained grader reviewed these images and identified the ones where the disc was present but incorrectly segmented, and these were included in the training set. In practice, these included many analog, macula-centered images from the Rotterdam sets, which specially challenged the model.
@@
For fovea detection, we used CFIs from a sub-cohort of the RS previously graded for signs of Age-related Macular Degeneration (AMD); which included localization of the fovea for ETDRS grid placement.
@@
Figure 1 shows the statistics of the final Rotterdam development sets. Vessel and artery-vein segmentation sets are similar (partly due to overlap) and are closer to the overall distribution of the RS. The disc segmentation dataset, on the other hand, contains more images from the analog devices, which were often the hardest to segment partly due to unconventional disc placement.
@@
Figure 1: Statistics of the Rotterdam development sets used to train and evaluate VascX models. NA indicates the information was not available.
CFI annotation
@@
Images were annotated using custom software for ophthalmological image annotation [33]. Four professional graders participated in the process. The graders worked on desktop computers with a drawing tablet for segmentation. The annotation process was different per annotation type:
@@
We followed an AI-assisted process by initializing the drawing interface with with outputs from vessel segmentation models. Initially (first 200 images) we made use of models trained on public data (including Automorph [57]) to later switch to a model trained on both public data and annotations from previous iterations of this process. The graders were instructed to independently fix the masks produced by the model, including missing vessels and over-segmentation. A digitally enhanced version of the image with increased contrast of vessel edges was used to aid the graders (Figure 2(a)).
@@
We resolved connectivity issues at artery-vein crossings by annotating arteries and veins on separate layers. An interface was developed specifically for artery-vein segmentation (Figure 2(b)) to allow the graders to start from a vessel segmentation (no A/V distinction) and color arteries and veins into two independent masks or layers using drawing tools. A third layer - Unknown was added for vessels that could not be recognized as either. They were asked to color A/V crossings on both layers (overlapping). They were able to visualize and correct each layer independently, or all the same time. Another feature allowed them to visualize the connected components of each mask in different colors to easily find mistakes in connectivity. The entire process consisted in 1) correcting mistakes in the provided mask (vessel segmentation); 2) coloring the mask into A/V and unknown colors/layers; 3) Verifying connected components on the A/V masks and filling any gaps.
@@
The fovea was localized by four graders who followed a standard grading protocol which included placing an ETDRS grid; overlaid on the images.
@@
Figure 2: Screen capture of the software used for all the annotations on Rotterdam CFIs.
Model Training
@@
    Detection of the CFI boundaries defined by a circle and top, bottom, left and right lines (see Figure 3).
@@
The geometric steps (cropping and resizing) were replicated on the labels for training. Figure 3 shows outputs of our pre-processing algorithm. Models were trained on the concatenation of the cropped CFI and its contrast-enhanced version (6 input channels).
Figure 3: Sample of pre-processed images from the Rotterdam sets, showing: original images (first row), the results of bounds detection defined by the intersection of a circle and optional lines (second row) and the cropped and contrast-enhanced image (third row).
Vessel, A/V and Disc Segmentation Models
@@
Vessel annotations were encoded as 0) background and 1) vessel. Artery-vein annotations were encoded as 0) background, 1) arteries, 2) veins, and 3) crossings (both artery and vein). We used this approach for A/V segmentation with the goal of boosting the relative importance of the crossings in the loss function. The loss function was the sum of Dice and Cross-entropy with equal weights. The optic disc segmentation model was similarly trained on binary masks encoded as 0) background and 1) disc mask; with the same loss function. Figure 6(a) shows sample training batches for the vessel and artery-vein segmentation models.
@@
Fovea Localization Model
The fovea localization model was trained separately to localize keypoints on the CFIs. A heatmap regression approach [55] with the same U-Net architecture used for artery-vein segmentation as backbone. Heatmap regression consists in training the model to regress a probability image or heatmap of the keypoint location. The pixel location of the maximum on the output heatmap is then taken as the output keypoint location. Train heatmaps may be generated using different probability density functions centered at the ground truth keypoint location. We generated target heatmaps ($512x512px$) using a Gaussian with constant $\sigma$ centered at the keypoint. The value of $\sigma=50px$ was selected via experiments on a holdout set for the task of localizing the disc center on CFIs. Mean squared error (MSE) was used as loss function to train the model to approximate the continuous target heatmaps. The location (2D-index) of the maximum over the heatmaps output by the model was taken as the keypoint location.
Input CFIs were augmented via random appearance transforms: de-focus, hue-saturation-value, Gamma and Gaussian noise; followed by geometric transforms: horizontal flipping, random scale (1.0-1.05 times original scale) random rotation (up to 5 degrees) and finally resizing down to model input size. Test-time augmentation was correspondingly applied using the vertical flip of the input. Figure 7 shows sample training batches for the disc segmentation and fovea localization models.
Quality Classification Model